sue can shovel snow from her driveway in 40 minutes. Jim can do the same job in 50 minutes. How long would it take Sue and Jim to shoved the driveway if they worked together?
Answer by Alan3354(69443)

sue can shovel snow from her driveway in 40 minutes. Jim can do the same job in 50 minutes. How long would it take Sue and Jim to shoved the driveway if they worked together?
--------------
In one minute, Sue shovels 1/40.
In one minute, Jim shovels 1/50.
Together, they to 1/40 + 1/50 per minute.
1/40 + 1/50 = 9/200 of the drive per minute.
The minutes per driveway = 200/9
--------------------------------
A quicker way to do these is the product/sum
40*50/(40+50) = 2000/90 = 200/9